# An alternative theory of magnetic flux tubes in strong fields via axion origin photons

### v2

@@ I Introduction
Figure 13: The time variations of total solar irradiance (TSI; orange curve; see Fig. 9 in (J. Pelt & O. Kärner, 2012)), simulated solar magnetic flux (black curve; see Fig. 7 in M. Dikpati et al. (2008); and also Fig. 16 in V. D. Rusov et al. (2013)) (both curves were smoothed with 29-year weighted moving average); the geomagnetic field secular variations (the Y-component; red curve) measured at the Eskdalemuir observatory (World Data Centre of Geomagnetism (2025)World Data Centre of Geomagnetism (Edinburgh), Edinburgh); variations in the Earth’s rotation speed (green curve; joined data from Fig. 3 in N. S. Sidorenkov (2016), Fig. 3.9 in N. S. Sidorenkov (2009), and the EOP C01 IAU1980 (1846-now) dataset by International Earth Rotation and Reference Systems Service (2025)); the variations of the average global ocean level (Pacific Decadal Oscillation (PDO), NOAA Physical Sciences Laboratory (2025) + Atlantic Multidecadal Oscillation (AMO), NOAA Physical Sciences Laboratory (2023); blue curve) with a trend in mm/years near 60-year oscillation (see Fig. 1 in D. P. Chambers et al. (2012)). The latter three curves were first smoothed using 5-year moving average, and then the 11-year moving average. All considered time variations between the Sun and the Earth always have a time interval of approximately $\sim 10\div 11$ years. The only serious disagreements between the reconstructions and observations occur during the Second World War, especially in the period 1944-1945 (see C. K. Folland et al. (2018)).
@@
First, we recall that the physics of experimental data related to the anticorrelation between 11-year variations in the luminosity of the Sun and the Earth (see Fig. 13) currently cannot explain one of the most mystical phenomena of solar-terrestrial physics, which is contained in well-known experiments, for example, between the temporal variations of the magnetic flux of the solar tachocline (black curve; see Fig. 7 in M. Dikpati et al. (2008); and also Fig. 16 in V. D. Rusov et al. (2013)), and secular variations in the Earth’s magnetic field (Y-component from World Data Centre of Geomagnetism (2025)World Data Centre of Geomagnetism (Edinburgh) (Edinburgh)), or variations in the Earth’s rotation speed (see Fig. 3 in N. S. Sidorenkov (2016), Fig. 3.9 in N. S. Sidorenkov (2009), and EOP C01 IAU1980 (1846-now) dataset by International Earth Rotation and Reference Systems Service (2025)), as well as the variations in the mean global sea level (Pacific Decadal Oscillation (PDO) (NOAA Physical Sciences Laboratory, 2025) + Atlantic Multidecadal Oscillation (AMO) (NOAA Physical Sciences Laboratory, 2023)) with a trend in mm/years near 60-year oscillation (see Fig. 1 in D. P. Chambers et al. (2012)).
From here follows the main argument for solving the above problem (see Fig. 13; and also Fig. 6 in V. D. Rusov et al. (2021)). We understand that it is the presence of asymmetric dark matter (ADM) in the Earth’s core that can, on the one hand, be the reason for solving the density deficit in the inner core, and on the other hand, achieve the necessary consensus (see D. Antonangeli et al. (2018); E. Edmund et al. (2019)) not only regarding the absolute values of the solid iron velocities under the conditions of the inner core, but also regarding the iron melt velocities from pressure and temperature. If this is true, then we can prove that the inner core is not only a hidden reservoir of hydrogen on Earth (see e.g. T. Sakamaki et al. (2016), but also the cause of the existence of a heat flow source from the Earth’s core, which, based on ADM variability (see V. D. Rusov et al. (2021)), generates 11-year thermal, magnetic and temperature cycles in all parts of the Earth, including from the Earth’s surface to the troposphere.
This means that when the maximum of magnetic fluxes on the Earth’s surface exceeds the energy of solar irradiation in the troposphere, and the minimum of these fluxes on the Earth’s surface becomes equal to solar irradiation in the troposphere, then the 11-year luminosity of the Earth will anticorrelate with the 11-year luminosity of the Sun.
@@ III Summary and Outlook
The main result is the following: three “free” parameters – the magnetic field in the tachocline of the order of $\sim 10^{7}~G$ (see Fig. (A.1) and Eq. (A17) in V. D. Rusov et al. (2021)), axions with mass $m\sim 3.2\cdot 10^{-2}~eV$ and coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$ (see Eq. (11) in V. D. Rusov et al. (2021)) and the asymmetry dark matter (ADM) in the Universe with $m\sim 5~GeV$ (see V. D. Rusov et al. (2021); A. C. Vincent et al. (2016)) – are the source of a complete solution of the problem of an alternative theory of magnetic flux tubes in strong fields using 11-year variations of axion origin photons, which are caused by anticorrelations of 11-year variations of ADM density, gravitationally captured on the Sun.
From here, we understand that the fraction of dark matter in the Universe, which is about $\sim$27%, should be the reason for the appearance of a very large amount of dark matter with $m\sim 5~GeV$, for example, in our Milky Way galaxy. This means that the dark matter ADM in our galaxy should be not only in stars, but also on planets, for example, the Sun and the Earth.
As not surprisingly, a simple but fundamental conclusion emerges. The anticorrelation of the 11-year variations in the ADM density, gravitationally trapped both on the Sun and on the Earth, must necessarily participate in all the physical problems of dark matter, which are related, for example, not only to our solution of magnetic flux tubes in strong fields, describing the physics of sunspots, but also to the solution of the global climate of the Earth, describing the physics of global warming due to temperature and magnetic fields, but, crucially, without the existence of anthropogenic enhancement of the greenhouse effect.
The latter suggests that, in an intriguing perspective, the experimental observation of anticorrelations between 11-year variations in the luminosity of the Sun and the Earth (see Fig. 13) leads to the emergence of dark matter asymmetry in the Earth’s core, which, as not surprisingly, controls, among other things, the magnetic field from the core to the Earth’s surface. This means that when the maximum magnetic field at the Earth’s surface exceeds the solar irradiation energy in the troposphere, and the minimum of such energy flux at the Earth’s surface is equal to the solar irradiation energy in the troposphere, then the 11-year luminosity of the Earth will anticorrelate with the 11-year luminosity of the Sun.
This opens up a powerful prospect for observing both the axions, which in this mass range may be the main component of dark matter, and the anticorrelation of 11-year variations in the density of ADMs gravitationally trapped on both the Sun and Earth.

### v3

@@ II.2.2 Double maxima of 11-year sunspot cycles and the Gnevyshev gap
Figure 10: The time variations of total solar irradiance (TSI; orange curve; see Fig. 9 in [125]), simulated solar magnetic flux (black curve; see Fig. 7 in [126]; and also Fig. 16 in [127]) (both curves were smoothed with 29-year weighted moving average); the geomagnetic field secular variations (the Y-component; red curve) measured at the Eskdalemuir observatory [128]; variations in the Earth’s rotation speed (green curve; joined data from Fig. 3 in [129], Fig. 3.9 in [130], and the EOP C01 IAU1980 (1846-now) dataset [131]); the variations of the average global ocean level (Pacific Decadal Oscillation (PDO) [132] + Atlantic Multidecadal Oscillation (AMO) [133]; blue curve) with a trend in mm/years near 60-year oscillation (see Fig. 1 in [134]). The latter three curves were first smoothed using 5-year moving average, and then the 11-year moving average. All considered time variations between the Sun and the Earth always have a time interval of approximately $\sim 10\div 11$ years. The only serious disagreements between the reconstructions and observations occur during the Second World War, especially in the period 1944-1945 (see [135]).
From here we now return to the physics of the Gnevyshev gap. The answer is quite simple. The Gnevyshev intervals in the butterfly diagrams are related to the 11-year modulations of the ADM density, trapped on the Sun (via vertical density waves from the disk to the solar neighborhood) and, as a result, to anticorrelated indicators of variations in the sunspot cycles or photons of axion origin. When the ADM density in the solar interior is low (or the sunspot count is high), then the Gnevyshev intervals will be short in time. Conversely, when the ADM density in the solar interior is high (or the sunspot count is low), then the Gnevyshev intervals will be longer in time. As no surprisingly, it is the presence of ADM density modulations that shows that the appearance of “our indicated” Gnevychev gap is the result of precisely 11-year variations of butterfly cycles with sunspots.
The most interesting thing is that it is the 11-year variations of sunspot cycles or photons of axion origin (see Figs. 5a and 6b in [14]) that are anticorrelated not only with the 11-year modulations of ADM density on the Sun, but also with the 11-year modulations of ADM density in the solar neighborhood, and, in particular, on Earth (see e.g. Fig. 10).
@@
First, we recall that the physics of experimental data related to the anticorrelation between 11-year variations in the luminosity of the Sun and the Earth (see Fig. 10) currently cannot explain one of the most mystical phenomena of solar-terrestrial physics, which is contained in well-known experiments, for example, between the temporal variations of the magnetic flux of the solar tachocline (black curve; see [126] and Fig. 16 in [127]), and secular variations in the Earth’s magnetic field (Y-component from [128]), or variations in the Earth’s rotation speed (see [129, 130] ), as well as the variations in the mean global sea level (Pacific Decadal Oscillation (PDO) [132] + Atlantic Multidecadal Oscillation (AMO) [133]) with a trend in mm/years near 60-year oscillation (see Fig. 1 in [134]), between which there is a time lag of about 10-11 years, respectively (see Fig. 10; and also Fig. 6 in [14].
@@ III Summary and Outlook
Most importantly, the existing void in the tube causes high velocities and short lifetimes of sunspots from the tachocline to the photosphere, which is consistent with experimental data. This means that the modernity of a well-developed helioseismological inversion, surprisingly, persists to this day.
The perspective is to understand the physics of the anticorrelation between 11-year variations of total solar irradiance and 11-year variations of the magnetic field and other Earth parameters (see Fig. 10).